OpenStudy (anonymous):

If a varies directly as b, and a = 28 when b = 7, find b when a = 5.

OpenStudy (anonymous):

ratio is 28/7 which is equal to 4/1 so now cross multiply: 4/1 = 5/b 1 * 5 = 5 4 * b = 5 4 * 1.25 = 5 so b = 1.25 when a = 5

OpenStudy (anonymous):

Hope this helps!